OpenStudy (anonymous):

Arcsec -(2sqrt3)/(3)

OpenStudy (jdoe0001):

\(\bf sec^{-1}\left(-\cfrac{2\sqrt{3}}{3}\right)\qquad recall\quad sec(\theta)=\cfrac{1}{cos(\theta)}\implies cos(\theta)=\cfrac{1}{sec(\theta)}\quad\\ thus\\ \quad \\ sec^{-1}\left(-\cfrac{2\sqrt{3}}{3}\right)\implies sec(\theta)=\left(-\cfrac{2\sqrt{3}}{3}\right)\implies \cfrac{1}{\left(-\frac{2\sqrt{3}}{3}\right)} = cos(\theta)\\ \quad \\ -\cfrac{3}{2\sqrt{3}} = cos(\theta)\implies -\cfrac{3}{2\sqrt{3}} \cdot \cfrac{\sqrt{3}}{\sqrt{3}}= cos(\theta)\implies -\cfrac{\cancel{3}\sqrt{3}}{\cancel{6}}= cos(\theta)\\ \quad \\ \cfrac{\sqrt{3}}{2}= cos(\theta)\)

OpenStudy (dumbcow):

\[\sec(x) = -\frac{2\sqrt{3}}{3}\] sec = 1/cos so 1/sec = cos \[\cos(x) = -\frac{3}{2\sqrt{3}} = -\frac{3\sqrt{3}}{6} = -\frac{\sqrt{3}}{2}\] using unit circle x = 150 degrees

OpenStudy (dumbcow):

lol

OpenStudy (jdoe0001):

negative for that matter \(\bf -\cfrac{\sqrt{3}}{2}= cos(\theta)\)